OpenStudy (anonymous):

In the transpose matrix, are the elements in a matrix already in order, or do I have to arrange them in order?

OpenStudy (zzr0ck3r):

\[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]^t=\left[\begin{matrix}a & c \\ b & d\end{matrix}\right]\]

OpenStudy (zzr0ck3r):

I'm not sure what you are asking but I hope this helps.

OpenStudy (anonymous):

what if the elements that I want to transpose are not in order do I have to arrange them in order, so that I can transpose them?

OpenStudy (zzr0ck3r):

order?

OpenStudy (mandre):

No. Just switch the rows and columns around like zzr0ck3r said in his first post.

OpenStudy (mandre):

Transpose basically means each row becomes a column and each column becomes a row.

OpenStudy (anonymous):

oh ok thank you!